the world speed record on water was set on october 8, 1978 by ken warby. if ken drove his motorboat a distance of 1000m in 7.045 s, how fast was his boat moving? in m/s and in miles/hour and kph.


1
Expert's answer
2021-09-09T10:50:02-0400

"S = 1000m"

"t = 7.045s"

"1miles = 1609.34m"

"1km = 1000m"

"1h = 3600s"

"S = \frac{1000}{1609.34}\approx0.62miles"


"t = \frac{7.045}{3600}\approx0.00196h"


"v = \frac{S}{t}"


"v = \frac{1000}{7.045}=141.94\frac{m}{s}"


"v = \frac{0.62}{0.00196}=316.32\frac{miles}{h}"


"v = \frac{1}{0.00196}=510.2\frac{km}{h}"


"\text{Answer: }141.94\frac{m}{s};316.32\frac{miles}{h};510.2\frac{km}{h}"